### v2

# Bremsstrahlung Gamma Ray as a Probe of Short-Range Correlations in Nuclei

Abstract
Atomic nuclei and dense nucleonic matter in neutron stars exhibit short-range correlations (SRCs), where nucleons form temporally correlated pair in close proximity. The impact of SRC on heavy-ion collisions (HICs) --which create dense nuclear matter in terrestrial laboratory-- has not been explored until now. In this letter, we present the neutron-proton bremsstrahlung $\gamma$-ray emission from HICs as a novel indicator of SRC in nuclei. By analyzing the hardening of the bremsstrahlung $\gamma$-ray spectrum, which results from the enhancement of high-momentum components above the Fermi surface in nucleon momentum distributions, we precisely determine the SRC fraction in nuclei to be $(20\pm 7)\%$ at $3\sigma$ confidence level. This experiment provides the first direct and accurate signature of SRC in low-energy HICs, offering a new method to study the parton distribution function of nucleons in nuclei.
Introduction -- Nucleonic matter, composed of strongly interacting protons and neutrons, exhibits complex many-body dynamics. Among these, some nucleons form short-range correlations (SRCs) [1, 2], which are temporal fluctuation effects that go beyond the mean-field approximation. SRCs have been experimentally observed and provide an explanation for the high-momentum tail (HMT) observed beyond the Fermi surface in nucleon momentum distributions [1, 3]. These high-momentum components significantly influence the properties of nuclear matter, particularly the density dependence of the nuclear symmetry energy [4, 5, 6], a quantity that is crucial in dense environments like neutron stars and heavy-ion collisions (HICs).
Studies indicate that SRCs are primarily characterized by neutron-proton ($np$) pairs in the isosinglet (T=0) channel at short distances, favoring $ud$ diquark isosinglet and spin-singlet configurations [7]. These structural features of SRCs have been shown to significantly influence the EMC effect [8] in nuclei [9, 10, 11], establishing a direct link between short-range nucleon correlations and quark-level modifications of nuclear structure [2, 12, 13, 14, 15, 16]. Furthermore, a recent global analysis based on the QCD factorization scheme has extracted the universal effective parton distribution within SRC pairs and the mean-field fraction using a variety of high-energy data [17]. These findings suggest that SRC studies may offer valuable insights into the modification of parton distribution functions (PDFs) in the nuclear environment.
The majority of our present understanding of SRC comes from the electron-nucleus scattering [18, 19, 20, 1, 3] and proton-nucleus knockout [21, 22, 23] experiment. Among various experimental approaches, HICs offer a unique opportunity to create bulk dense nucleonic matter in terrestrial laboratory. Therefore, searching for signatures of SRC in HICs is of importance, as it not only links SRC to dense nucleonic matter, but also opens new avenues for quantifying the SRC fraction under extreme conditions. When two heavy nuclei collide, two-body $np$ collisions generate bremsstrahlung $\gamma$-rays [24, 25, 26]. Due to the influence of the high momentum tail (HMT) arising from SRCs, the average nucleon kinetic energies of both projectile and target increase significantly, resulting in a hardening of the emitted $\gamma$-ray spectrum, particularly at high energies [27]. In comparison to the hadron probes [28, 29, 30], the high-energy $\gamma$-rays are largely unaffected by the nuclear medium, making them a clean and unambiguous probe for detecting SRC effects [27, 31, 32].
Theoretical calculations using the Isospin-dependent Boltzmann-Uehling-Uhlenbeck (IBUU) transport model [33], incorporating Momentum-Dependent Interaction (MDI) [34], predict the bremsstrahlung $\gamma$-ray spectra produced in HICs by varying the HMT fraction ($R_{\text{HMT}}$), which quantifies the proportion of nucleons with momenta above the Fermi surface in nuclei [27]. By comparing model predictions with experimental $\gamma$-ray spectrum, The HMT value $R_{\text{HMT}}$ can be effectively constrained. In our previous experiment involving $\rm{}^{86}Kr$+$\rm{}^{124}Sn$ collisions, the bremsstrahlung $\gamma$-ray spectrum was measured, delivering a nonzero HMT fraction induced by SRCs [35, 36]. The results showcased the potential of bremsstrahlung $\gamma$-ray emission as a novel probe of SRC in nuclei.
This letter reports on a new measurement of bremsstrahlung $\gamma$-rays from $\rm{}^{124}Sn$+$\rm{}^{124}Sn$  collisions at 25 MeV/u with significantly improved statistics and extended energy range. The $\gamma$-ray spectrum is analyzed using the IBUU-MDI transport model, which incorporates the $np\rightarrow np\gamma$ process. The SRC-induced HMT is included in the initialization of the nucleon momentum. A precise $R_{\text{HMT}}$ is obtained, providing a new paradigm to study SRC of nucleons and their underlying partonic structure in nuclei through HICs.
Experiment-- The experiment was performed on Radioactive Ion Beam Line 1 (RIBLL1) at the Heavy Ion Research Facility in Lanzhou (HIRFL). The products from the reaction of 25 MeV/u $\rm{}^{124}Sn$+$\rm{}^{124}Sn$  were measured by the Compact Spectrometer for Heavy Ion Experiment (CSHINE) [37, 38, 39]. Fig. 1 presents the detector setup. The charged isotopes of $Z\leq 5$ were measured by 6 silicon-strip detector telescopes (SSDTs) covering the polar angle $20^{\circ}<\theta_{\rm lab}<60^{\circ}$ in partial azimuth [40]. Neutrons are detected by a $4\times 5$ array of plastic scintillator installed at 2.0 meters to the target. With 4 fast $\rm BaF_{2}$ $\gamma$-detectors surrounding the target delivering starting time (T0) information, neutrons can be discriminated from the $\gamma$-rays [37, 41]. Besides, a liquid scintillator (LS) neutron detector in size of $\rm\phi 348~{}mm\times 1020~{}mm$ is mounted at 3 m to the target (not shown in Fig. 1).
Figure 1: An illustration of the CSHINE detector setup. A schematic view of an $\rm{}^{124}Sn$+$\rm{}^{124}Sn$  event is plotted, involving the $np$ collision that emits a bremsstrahlung $\gamma$-ray, subsequently detected by the CsI(Tl) hodoscope with three surrounding plastic scintillators.
To measure the energetic bremsstrahlung $\gamma$-rays from the reactions on the target, a $4\times 4$ CsI(Tl)-based hodoscope has been mounted [42]. Each unit has a dimension of $70\times 70\times 250~{}\rm mm^{3}$. Three plastic scintillators were installed surrounding the CsI(Tl) hodoscope to veto the cosmic ray background as well as the $\gamma$ events with energy leaking to the outside. A dual-range amplifier was applied to record the small and large energy signals, and then the response of each CsI(Tl) unit was calibrated by a dual-range method [43]. The energy range of each individual CsI(Tl) unit was optimized in order to reach the maximum $\gamma$-ray energy deposit of 100 MeV in the hodoscope. To subtract the background, which mainly includes cosmic-ray muons and residual $\gamma$ radiation at low energies, the background spectrum was measured when the beam was off.
The data acquisition was triggered by one of six conditions listed below. $\rm Trg_{1}$, one CsI (Tl) unit and one SSDT fired (SSD M1 & CsI M1); $\rm Trg_{2}$, two SSDT fired (SSD M2); $\rm Trg_{3}$, one neutron array unit and one SSDT fired (SSD M1 & NA M1); $\rm Trg_{4}$, liquid scintillator and one SSDT fired (SSD M1 & LS); $\rm Trg_{5}$, one neutron array unit and one T0 fired (NA M1 & T0); $\rm Trg_{6}$, liquid scintillator and one T0 fired (LS & T0). The trigger scheme was implemented using field-programmable gate array (FPGA) technology [44].
Figure 2: Event-wise vertical spreading $\delta_{y}$ as a function of the total reconstructed energy $E_{\rm tot}$ for beam-on (a) and beam-off (b) data. Correlation between the energy deposited in the event core $E_{\rm core}$ and $E_{\rm tot}$ for beam-on (c) and beam-off (d) data. A typical display of a high energy $\gamma$ event from the reaction (e), and of a cosmic muon penetration event through the $\gamma$-hodoscope (f).
After calibrating the CsI(Tl) hodoscope, the energy deposits are reconstructed event-by-event. In each triggered event, the reconstruction algorithm identifies the clustered CsI(Tl) hits coincided in a $|\Delta T|<50$ ns time window and led by a ‘core’ with the largest energy deposit. The $\gamma$ events and the penetrating cosmic muon events can be distinguished by their topological characteristics, as depicted in Fig. 2 in beam-on (left panels) and beam-off (right panels) stages, respectively. For the beam-on measurement, on the scattering plots of the variance of the spatial distribution of the fired units $\delta_{y}$ [42] (a) and the energy deposit in the core $E_{\rm core}$ (c) as a function of the total reconstructed energy $E_{\rm tot}$, two dominant components are clearly visible. One is the $\gamma$ event of small spatial extension and in the vicinity of $E_{\rm core}\approx E_{\rm tot}$, indicating that most of the energy is collected by the event core with less extension to the neighbouring units. The other is dominated by the cosmic muon characterized by large $\delta_{y}$ and $E_{\rm tot}>E_{\rm core}$, indicating the incident muon fires several CsI(Tl) units with approximate energy deposit. The events with large $E_{\rm tot}$ in the beam-on data exhibit similarity with those observed in the beam-off background, as shown in (b) and (d), where cosmic muons dominate in large $E_{\rm tot}$ area. Fig. 2(e) and (f) illustrate a typical $\gamma$-ray event and a penetrating cosmic muon event, respectively. Low-energy events ($E_{\rm tot}<30$ MeV) with multiple hits mainly result from statistical residual $\gamma$ emissions.
Fig. 3 presents the total energy spectra in beam-on (black) and beam-off (red) measurements. No selection on the impact parameter $b$ is done, since the spectrum shape shows insignificant dependence on $b$ [35]. The beam-off background is scaled to match the beam-on spectrum using the counts above $110~{}\rm MeV$. The pure $\gamma$ ray spectrum in laboratory can be obtained by subtracting the scaled background spectrum from the total spectrum, as displayed in the inset. It is noted here that the spectrum is affected by detector efficiency, and hence the theoretical spectra must be filtered by the same detector response before comparison can be made (see next). The detector filter matrix, describing the response of the CsI(Tl) hodoscope to $\gamma$-rays of varying energies in laboratory, was generated using Geant4 simulations [42].
Figure 3: The experimental $\gamma$ energy spectrum in laboratory measured with beam-on (black), compared to the background spectrum with beam-off (red), normalized to the counts above $110~{}\rm MeV$. The inset displays the background-subtracted $\gamma$ spectrum in laboratory.
The main systematic uncertainty originates from detector calibration with the linear response assumption, as the radioactive sources cover energies below 3 MeV, whereas the bremsstrahlung $\gamma$-rays may deposit tens of MeV in a single CsI(Tl) unit. Higher energy $\gamma$-ray beam test was conducted to characterize the nonlinearity of the CsI(Tl) unit. As a result, the detector response correction functions (DRCFs) were obtained to quantify the deviations from linear calibration. The uncertainty brought by the linear calibration is within 4% [43].
Results and Discussions -- The IBUU transport model [45, 46, 47] simulates nucleus-nucleus collisions and has been extended to include isospin degrees of freedom [48] and momentum-dependent nucleon interaction potential (MDI) [49]. The SRC and HMT effects are introduced in the initialization of single-nucleon momentum distribution $\rho(k)$ [50, 27]. For the asymmetric nuclear matter (ANM), the high-momentum component depends linearly on the isospin asymmetry $\delta$ [51, 52, 6]. Thus $\rho(k)$ for ANM is parameterized as
$\rho_{\rm N}(k)=\left\{\begin{array}[]{rcl}B_{\rm N},&&{k\leq k^{\rm N}_{\rm F}}\\ \eta(1+s\delta)/k^{4},&&{k^{\rm N}_{\rm F}<k\leq\lambda k^{\rm N}_{\rm F}}\\ 0,&&{k>\lambda k^{\rm N}_{\rm F}}\end{array}\right.$ (1)
[…]
Due to the low production rate of the bremsstrahlung $\gamma$, its impact on the nucleon kinematics is negligible. Perturbation method is then employed to calculate the photon production probability in each $np$ collision. Based on the one boson exchange (OBE) model, a well fitted formula is applied to calculate the elementary double differential $\gamma$ production probability [26],
$\frac{d^{2}P}{d\Omega dE_{\gamma}}=1.67\times 10^{-7}\times\frac{[1-(E_{\gamma}/E_{\rm max})^{2}]^{\alpha}}{E_{\gamma}/E_{\rm max}}.$ (2)
Here, $E_{\gamma}$ represents the energy of produced photons, and $E_{\rm max}$ is the total available energy in the center-of-mass (CM) frame of the colliding $np$ pair. The coefficient $\alpha$ reads $\alpha=0.7319-0.5898\beta_{i}$, where $\beta_{i}$ is the nucleon velocity. The total $\gamma$ yield per event is obtained by summing over all photon-producing $np$ collisions during the reaction. The $\gamma$ spectrum show minimal dependence on the nuclear mean-field, impact parameter, and symmetry energy in the test, as expected because bremsstrahlung photons are predominantly emitted in the early stages and interact only electromagnetically.
Fig. 4 (a) illustrates the rebinned experimental $\gamma$ spectrum converted to the CM frame. The black dots indicate the central values with statistical errors, while gray bands represent systematic uncertainties. Several key theoretical curves, filtered by the detector response in laboratory, are presented for comparison. Unlike our previous experiment [35, 36], the $\gamma$ detector array was operated in active triggering mode, preventing the determination of the number of $np$ collisions in HICs for normalizing the $\gamma$ spectrum. Therefore, the comparisons with model predictions focus on spectral shape rather than absolute yield. Theoretical spectra were uniformly scaled by $2.5\times 10^{9}$ to match the data range, without affecting $R_{\rm HMT}$ extraction based on shape analysis.
Figure 4: (a) Comparison of the rebinned experimental $\gamma$ spectrum (black dots) in CM reference with various detector-filtered theoretical curves. Statistical (systematic) uncertainties are denoted by vertical bars (gray bands). (b) Likelihood distribution as a function of $R_{\rm HMT}$ with quadratic fit (dashed curve). The error bars represent the standard deviation of likelihood values by varying the calibration and detector response data sets. Confidence levels of $3\sigma$ and $5\sigma$ are indicated, respectively.
A maximum likelihood analysis was performed to compare the model prediction with the experimental data, using the likelihood function [35] defined as
$\ln L^{\prime}(R_{\text{HMT}})=\sum_{i}^{\text{range}}n_{i}\ln p_{i}(R_{\text{HMT}}).$ (3)
where $i$ sums over the experimental points within the statistical analysis interval, and $n_{i}$ denotes the count in the $i^{\rm th}$ experimental bin. The quantity $p_{i}$ represents the probability that the theoretical model predictions fall into the corresponding histogram bin under the specified statistical analysis interval for a given $R_{\text{HMT}}$. $\gamma$ rays below $30~{}\rm MeV$ are dominated by collective resonances, statistical emissions and possibly residual radiation in the target chamber. Accordingly, the analysis was restricted to the range of interest $35<E_{\gamma}^{\rm cm}<100~{}\rm MeV$ to evaluate the likelihood between experimental spectra and IBUU-MDI predictions for different $R_{\rm HMT}$ values.
[…]
The result reveals a new paradigm to probe the PDF of nucleon in nuclei (nPDF) via low energy HICs. According to the collinear factorization scheme [54], the nPDF parametrization is composed of a linear combination of PDFs of free nucleons and of SRC pairs, with the fraction fully encapsulated in the ratio of neutrons (protons) in SRC pairs $C_{n}^{\rm A}$ ($C_{p}^{\rm A}$). With $R_{\text{HMT}}=20\%$ in 124Sn ($A=124$, $\delta=0.194$), one writes $C_{n}^{124}=16\%$ and $C_{p}^{124}=24\%$, respectively, in accordance with the results obtained from high energy data [54]. The result also underscores the need for further systematic experimental studies of higher-energy $\gamma$-ray production in HICs. Interestingly, based on very similar mechanism of the presence of HMT, sub-threshold particle production in proton-nucleus collisions is proposed recently as a probe of SRC in nuclei [55].
Summary -- The precise measurement of the $np$ bremsstrahlung $\gamma$-rays in $\rm{}^{124}Sn$+$\rm{}^{124}Sn$  at 25 MeV/u not only resolves a longstanding puzzle that emerged in the 1980s regarding the origin of high-energy $\gamma$-rays in HICs [56], but also introduces a new indicator of SRC, a significant phenomenon that is crucial to understand nuclear structure and the properties of dense nucleonic matter. By exploiting the advantage of HICs, where the $np$ bremsstrahlung $\gamma$-ray energy spectrum is hardened due to the HMT of nucleons resulting from SRC, we compare the experimental $\gamma$ spectrum with IBUU-MDI model predictions. This comparison allows us to precisely determine the high-momentum nucleon fraction in the 124Sn nucleus, $R_{\rm HMT}=(20\pm 7)\%$ at $3\sigma$ confidence level. Our experiment provides the first direct signature of SRC in low-energy HICs, establishing a new approach to study the parton momentum distribution of nucleons in nuclei.
Acknowledgements. This work is supported by the Ministry of Science and Technology under Grant No. 2022YFE0103400, by the National Natural Science Foundation of China under Grant Nos. 12335008, 12205160, 12275129, by the Center for High Performance Computing and Initiative Scientific Research Program in Tsinghua University and by HIRFL. The authors acknowledge the HIRFL machine team for providing the excellent beam condition.